OpenStudy (anonymous):

Caty has a job at a local library. She earns $7.25 an hour. She worked 20 hours last week, and she earns an extra $10.00 per week if she is on time every day. Caty also earns a bonus of 15% of her hourly daily pay if she sweeps the floor every night. How much did Caty earn for the past week of work if she was on time very day and swept the floor every night?

OpenStudy (anonymous):

ok so break it down. she earns 7.25/hour so that's 7.25 * 20 = 145. she earns an extra 10 / week if she's on time. she did that so that's 145 + 10 = 155. she also got an extra 15% /hour for sweeping. so (145*.15) = 21.75 so she made 145+10+21.75 = 176.75